OpenStudy (anonymous):

FAN + MEDAL FOR HELP!!! After the WAVA Pacific Science Center outing, Brittany stops to get her pictures printed. The cost to print her pictures is $2.00 per order plus $1.00 for each print after the first print. Objective: Write a simplified equation to represent the cost c of printing p pictures.

OpenStudy (anonymous):

srry did that worng

OpenStudy (anonymous):

c=2+1p

OpenStudy (anonymous):

thats the equation

OpenStudy (dtan5457):

@cody77445 are you sure?

OpenStudy (anonymous):

that seems right to me am i wrong if so please correct

OpenStudy (anonymous):

i want people to get the right answer

OpenStudy (dtan5457):

The wording is a bit confusing here. "The cost to print her pictures is $2.00 per order plus $1.00 for each print after the first print." Who knows, I may be wrong too.. but reading this is telling me that the first print is $2, but any other print after that is $1. Would it not be the function 2+(x-1)? Because if it's just one print 2+(1-1)=$2<<<as the situation states the first print is $2 2nd print after that 2+(2-1)=$3<<<<<1st print is $2, and the 2nd one is $1,=$3

OpenStudy (anonymous):

there is 12 pictures to be printed srry i forgot to add tht